A boat sails 30km at a uniform rate. If the rate had been 1 km/h more, the time of the sailing would have been 1h less. Find the rate of travel.
Answer by stanbon(75887)

A boat sails 30km at a uniform rate. If the rate had been 1 km/h more, the time of the sailing would have been 1h less. Find the rate of travel.
-----------------
slower rate DATA:
dist = 30 km ; rate = x kph ; time = d/r = 30/x hrs
------
faster rate DATA:
dist = 30 km ; rate = (x+1) kph ; time = d/r = 30/(x+1) hrs
------
slower - faster = 1 hr
------
30/x - 30/(x+1) = 1
--------------------------------
30(x+1) - 30x = x^2+x
----------------------------------
x^2 + x - 30 = 0
------
(x+6)(x-5) = 0
-----
Ans: slower rate = x = 5 km/h
faster rate = x+1 = 6 km/h
----------------
Cheers,
Stan H.
------------